OpenStudy (anonymous):

which is equivalent to 3√16 + √54? a. 2 3√5 b. 5 3√2 c. 3√70 d. 6 √70

OpenStudy (solomonzelman):

think how to brake up the square root into the product f a bunch of perfect squares, like \[\sqrt{48}=\sqrt{16} \times \sqrt{3}~~~~or~~~~4\sqrt{3}\]

OpenStudy (anonymous):

how do u do it

OpenStudy (solomonzelman):

When you have\[3\sqrt{16}\]it's same as \[3\sqrt{4^2}=?~~~~~~~you~~tell~~me.\]

OpenStudy (anonymous):

i dont understand

OpenStudy (solomonzelman):

do you see that 16 is a perfect square?

OpenStudy (anonymous):

yea

OpenStudy (solomonzelman):

\[16=4^2\]

OpenStudy (anonymous):

but how do u get the answer

OpenStudy (solomonzelman):

\[So,~~~3\sqrt{16}=3\sqrt{4^2}=3 \times 4 = 12\]\[\sqrt{54}= \sqrt{9} \times \sqrt {6} = \sqrt{3^2} \times \sqrt{6}=3 \times \sqrt{6}=3\sqrt{6} \]\[SO~~~~~~~12+3\sqrt{6}\]

OpenStudy (anonymous):

theres answer choices

OpenStudy (solomonzelman):

The answer choices are wrong, or you asked the question incorrectly.

OpenStudy (anonymous):

i asked it right

OpenStudy (solomonzelman):

\[I~~~am~~~positive~~~that~~~I~~~am~~~ri ght!\]

OpenStudy (solomonzelman):

it can't be any of the given choices for the given question.